# A survey and benchmark of high-dimensional Bayesian optimization of discrete sequences

### v1

@@ Abstract
Optimizing discrete black-box functions is key in several domains, e.g. protein engineering and drug design. Due to the lack of gradient information and the need for sample efficiency, Bayesian optimization is an ideal candidate for these tasks. Several methods for high-dimensional continuous and categorical Bayesian optimization have been proposed recently. However, our survey of the field reveals highly heterogeneous experimental set-ups across methods and technical barriers for the replicability and application of published algorithms to real-world tasks. To address these issues, we develop a unified framework to test a vast array of high-dimensional Bayesian optimization methods and a collection of standardized black-box functions representing real-world application domains in chemistry and biology. These two components of the benchmark are each supported by flexible, scalable, and easily extendable software libraries (poli and poli-baselines), allowing practitioners to readily incorporate new optimization objectives or discrete optimizers. Project website: https://machinelearninglifescience.github.io/hdbo_benchmark.
@@ 1 Introduction
Optimizing an unknown and expensive-to-evaluate function is a frequent problem across disciplines (Shahriari et al.,, 2016): examples are finding the right parameters for machine learning models or simulators, drug discovery or train scheduling, to name a few. In some scenarios, evaluating the black-box involves an expensive process (e.g. training a large model, or running a physical simulation); Bayesian Optimization (BO, Močkus, (1975)) is a powerful method for sample efficient black-box optimization. High dimensional (discrete) problems have long been identified as a key challenge for Bayesian optimization algorithms (Wang et al.,, 2013; Snoek et al.,, 2012) given that they tend to scale poorly with both dataset size and dimensionality of the input.
High-dimensional BO has been the focus of an entire research field (see Fig. 1), in which methods are extended to address the curse of dimensionality and its consequences (Binois and Wycoff,, 2022; Santoni et al.,, 2023). Within this setting, discrete sequence optimization has received particular focus, due to its applicability in the optimization of molecules and proteins. However, prior work often focuses on sequence lengths and number of categories below the hundreds (see Fig. 2), making it difficult for practitioners to judge expected performance on real-world problems in these domains. We contribute (i) a survey of the field while focusing on the real-world applications of high-dimensional discrete sequences, (ii) a benchmark several optimizers in established black boxes, and (iii) an open source, unified interface: poli and poli-baselines.111https://github.com/MachineLearningLifeScience/{poli, poli-baselines}
@@ 2 Preliminaries
Bayesian optimization requires a surrogate model and an acquisition function (Garnett,, 2023). Given both, the objective function is sequentially optimized by fitting a model to the given observations and numerically optimizing the acquisition function with respect to the model to select the next configuration for evaluation. Frequently, the model is a Gaussian process (GP, Rasmussen and Williams, (2006)), and popular choices for the acquisition function are Expected Improvement (Jones et al.,, 1998) and the Upper Confidence Bound (Srinivas et al.,, 2012). A GP allows to express a prior belief over functions. Formally, it is a collection of random variables, such that every finite subset follows a multivariate normal distribution, described by a mean function , and a positive definite covariance function (kernel) (Rasmussen and Williams,, 2006, p. 13). Assuming that observations of the function are distorted by Gaussian noise, the posterior over the function conditioned on these observations is again Gaussian. The prediction equations have a closed form and can be evaluated in time where is the number of observations.
@@
Most HDBO methods are tested on toy examples, hyperparameter tuning, or reinforcement learning tasks (Binois and Wycoff,, 2022; Penubothula et al.,, 2021). We focus on discrete sequence optimization (DSOpt), which has several applications beyond the usual examples (e.g. MaxSAT, or Pest Control) (Papenmeier et al.,, 2024), and is key in applications to biology and bioinformatics (Gómez-Bombarelli et al.,, 2018; Stanton et al.,, 2022; Gruver et al.,, 2023). Drug design and protein engineering can be thought of as DSOpt problems, if we consider the SMILES/SELFIES representation of small molecules (Weininger,, 1988; Krenn et al.,, 2020), or the amino acid sequence representation of proteins (Needleman and Wunsch,, 1970).
@@
Binois and Wycoff, (2022) initially surveyed the field of high-dimensional GPs, focusing on applications to BO, and proposed a taxonomy of structural assumptions for GPs that includes variable selection, additive models, linear, and non-linear embeddings. This work has since been updated by Wang et al., (2023) and Santoni et al., (2023). The latter presents an empirical study on the continuous, toy-problem setting up to 60 dimensions and refines the taxonomy (Binois and Wycoff,, 2022) into five categories, separating trust regions from the rest. Our work is most similar to Dreczkowski et al., (2023)’s comprehensive overview of discrete BO (MCBO), and (Gao et al.,, 2022)’s benchmark of small molecule optimization. In both, HDBO is not in focus.
@@ 3 A taxonomy of high-dimensional Bayesian Optimization
We describe the field of high dimensional BO and the large number of related publications through a refined taxonomy building on previous work, discussing variable selection, additive models, trust regions, linear embeddings, non-linear embeddings, gradient information, structured spaces, and others in turn. While encompassing taxonomies over fields may initially appear ill-advised (Wilkins,, 1668, pp.22), we highlight commonalities in strategies that give structure to the HBDO problem-space.
We expand previous surveys (Binois and Wycoff,, 2022; Santoni et al.,, 2023) and identify a finer taxonomy, of seven method groups and new families of structured spaces (i.e. methods that work directly on mixed representations, or Riemannian manifolds, previously categorized as non-linear embeddings), and methods that rely on predicted gradient information. This new separation emphasizes the heterogeneous nature of discrete solvers: some optimizers work directly on discrete space (structured spaces), while others optimize using latent representations (non-linear embeddings); gradient-based methods are separated to show alternatives when first-order information is available. Fig. 1 presents a timeline of HDBO methods, split into these families, and all methods are detailed in supplementary Table LABEL:tab:appendix:full_taxonomy; methods are grouped according to their most dominant feature.
@@ 3.1 Variable selection
To solve a high-dimensional problem, one approach is to focus on a subset of variables of high interest.222Under the assumption that there exists an axis-aligned lower-dimensional active subspace. One selects the variables either by using domain expertise, or by Automatic Relevance Detection (ARD) (Rasmussen and Williams,, 2006, pp.106-107) i.e. large lengthscales indicate independence under the covariance matrix for GPs. Examples of this approach include Hierarchical Diagonal Sampling (HDS) (Chen et al., 2012b, ) and the Dimension Scheduling Algorithm (DSA) (Ulmasov et al.,, 2016). The former determines the active variables by a binary tree of subsets of , and fits GPs in lower-dimensional projections. DSA constructs a probability distribution by the principal directions of the training inputs and subsamples the dimensions accordingly. In contrast Li et al., (2018) randomly sample subsets of active dimensions.
@@ 3.2 Additive models
Han et al., 2021a select the decomposition which maximizes the marginal likelihood from a collection of randomly sampled decompositions, updating it every certain number of iterations. Alternatives include: leveraging a generalization based on restricted projections (Li et al.,, 2016), discovering the additive structure using model selection and Markov Chain Monte Carlo (Gardner et al.,, 2017), considering overlapping groups (Rolland et al.,, 2018), ensembles of Mondrian space-tiling trees (Wang et al.,, 2018), or use random tree-based decompositions (Ziomek and Bou Ammar,, 2023).
@@ 3.4 Linear embeddings
Instead of optimizing directly in input space , several methods rely on optimizing in a lower-dimensional space , which is linearly embedded into data space using a linear transformation (Wang et al.,, 2016). The matrix can be either selected at random (Wang et al.,, 2016; Qian et al.,, 2016), computed as a low-rank approximation of the input data matrix (Djolonga et al.,, 2013; Zhang et al.,, 2019; Raponi et al.,, 2020), constructed using gradient information and active subspaces (Palar and Shimoyama,, 2017; Nathan Wycoff and Wild,, 2021), or through the minimization of variance estimates (Hu et al.,, 2024).
@@ 3.5 Non-linear embeddings
Gómez-Bombarelli et al., (2018) pioneered latent space optimization (LSBO) by learning a latent space of small molecules through their SMILES representation using a Variational Autoencoder (VAE, Kingma and Welling, (2014); Rezende et al., (2014)), and optimizing metrics such as the qualitative estimate of druglikeness (QED) therein. Several approaches have followed, including smart retraining schemes that focus on promising points (Tripp et al.,, 2020), metric-learning approaches that match promising points together (Grosnit et al.,, 2021), and jointly learning the surrogate model and the latent representation (Maus et al.,, 2022; Lee et al.,, 2023; Chen et al.,, 2024; Kong et al.,, 2024). Stanton et al., (2022) take this further by learning multiple representations: one shared and required for both the decoder and surrogate, and one discriminative encoding as input for a GP used in the acquisition function. A prerequisite for these methods is a large dataset of unsupervised inputs, which may not be available in all applications. The methods that rely on training both the representation and the regression at the same time need supervised labels, which may be potentially unavailable. Optimization in embedding spaces greatly increases the complexity of problems that can be tackled, making it an appealing alternative for discrete sequence optimization in real-world tasks (see Fig. 2).
@@ 3.7 Structured spaces
Some applications work over structured spaces. For example, the angles of robot arms and protein backbones map to Riemannian manifolds (Jaquier et al., 2020a, ; Penner,, 2022), and input spaces might also contain mixed variables (i.e. products of real and continuous spaces). To compute non-linear embeddings (see Sec. 3.5) followed by standard Bayesian Optimization (or small variants thereof) can allow us to work over such spaces. Jaquier et al., 2020b use kernels defined on Riemannian manifolds (Feragen et al.,, 2015; Borovitskiy et al.,, 2020) and optimize the acquisition function using tools from Riemannian optimization (Boumal,, 2023). The authors expand their framework to high-dimensional manifolds by projecting to lower-dimensional submanifolds, which is roughly the equivalent to linear embeddings in the Riemannian settings (Jaquier and Rozo,, 2020).
@@ 4 Benchmarking the performance of HDBO methods
Practitioners that decide on what Bayesian optimization algorithm to use for their application will face several challenges. While surveying the field, we noticed two key discrepancies in the reported experimental set-ups: (i) the initialization varies from as low as none to five randomly/SOBOL sampled points to over , (ii) evaluation budgets also vary for the same types of tasks. Fig. 3 visualizes these different experimental set-ups as swarmplots. Moreover, our survey covered code availability. The state-of-the-art is being pulled by workhorses, which have democratized access to GP and BO implementations: GPyTorch (Gardner et al.,, 2018) and BoTorch (Balandat et al.,, 2020), and GPFlow (Matthews et al.,, 2017; van der Wilk et al.,, 2020) and Trieste (Picheny et al.,, 2023). These libraries are highly useful and impactful, yet one can obtain cross-dependency conflicts between them especially if third-party dependencies are introduced or if very specific versions are required for solver setups. As a particular example, solvers like ProbRep cannot co-exist with Ax-based solvers like SAASBO or Hvarfner’s VanillaBO. There is a need for isolating optimizers, specifying up-to-date environments in which they can run. These issues led to the development of poli.
@@
We want to solve truly high dimensional problems that are relevant for domains like biology and chemistry. To make the outcomes comparable, we require a unified way of defining the problem which includes consistent starting points, budgets, runtime environments, relevant assets (i.e.models used for the black-box), and a logging backend invoked for every oracle observation. To that end, we provide the Protein Objectives Library (poli) as a means to provide isolated black-box functions. Building on open source tools, poli currently provides 35 black-box tasks; besides the Practical Molecular Optimization (PMO) benchmark (Huang et al.,, 2021; Gao et al.,, 2022; Brown et al.,, 2019), it includes Dockstring (García-Ortegón et al.,, 2022) as well as other protein-related black boxes like stability and solvent accessibility (Delgado et al.,, 2019; Blaabjerg et al.,, 2023; Chapman and Chang,, 2000; Stanton et al.,, 2022). The majority of black-box functions can be queried with any string that complies with the corresponding alphabet making the oracles available for free-form optimization. This is an important distinction compared to pre-existing benchmarks that rely on pools of precompiled observations (Notin et al.,, 2023). We further provide an interface for the solvers used for the individual optimization tasks: poli-baselines. Consistent, stable (and up to date) environments of individual optimizers can be found therein, as well as a standardized way to query them and solve the problems raised in the previous section. These environments and optimizers are tested weekly through GitHub actions, guaranteeing their usability. Sec. A.4 provides a broader introduction to this software’s technical details.333Documentation: https://machinelearninglifescience.github.io/poli-docs/
4.2 Benchmarking HDBO on PMO
| HillClimbing | Hvarfner’s VanillaBO | RandomLineBO | BAxUS | Turbo | Bounce | ProbRep | |
| Oracle () | |||||||
| albuterol_similarity | |||||||
| amlodipine_mpo | |||||||
| celecoxib_rediscovery | |||||||
| deco_hop | |||||||
| drd2_docking | |||||||
| fexofenadine_mpo | |||||||
| gsk3_beta | |||||||
| isomer_c7h8n2o2 | |||||||
| isomer_c9h10n2o2pf2cl | |||||||
| jnk3 | |||||||
| median_1 | |||||||
| median_2 | |||||||
| mestranol_similarity | |||||||
| osimetrinib_mpo | |||||||
| perindopril_mpo | |||||||
| ranolazine_mpo | |||||||
| rdkit_logp | |||||||
| rdkit_qed | |||||||
| sa_tdc | |||||||
| scaffold_hop | |||||||
| sitagliptin_mpo | |||||||
| thiothixene_rediscovery | |||||||
| troglitazone_rediscovery | |||||||
| valsartan_smarts | |||||||
| zaleplon_mpo | |||||||

### v2

@@ Abstract
Optimizing discrete black box functions is key in several domains, e.g. protein engineering and drug design. Due to the lack of gradient information and the need for sample efficiency, Bayesian optimization is an ideal candidate for these tasks. Several methods for high-dimensional continuous and categorical Bayesian optimization have been proposed recently. However, our survey of the field reveals highly heterogeneous experimental set-ups across methods and technical barriers for the replicability and application of published algorithms to real-world tasks. To address these issues, we develop a unified framework to test a vast array of high-dimensional Bayesian optimization methods and a collection of standardized black box functions representing real-world application domains in chemistry and biology. These two components of the benchmark are each supported by flexible, scalable, and easily extendable software libraries (poli and poli-baselines), allowing practitioners to readily incorporate new optimization objectives or discrete optimizers. Project website: https://machinelearninglifescience.github.io/hdbo_benchmark.
@@ 1 Introduction
Optimizing an unknown and expensive-to-evaluate function is a frequent problem across disciplines (Shahriari et al.,, 2016), examples are finding the right parameters for machine learning models or simulators, drug discovery (Gómez-Bombarelli et al.,, 2018; Griffiths and Hernández-Lobato,, 2020; Pyzer-Knapp,, 2018), protein design (Stanton et al.,, 2022; Gruver et al.,, 2023), hyperparameter tuning in Machine Learning (Snoek et al.,, 2012; Turner et al.,, 2021) and train scheduling. In some scenarios, evaluating the black box involves an expensive process (e.g. training a large model, or running a physical simulation); Bayesian Optimization (BO, Močkus, (1975)) is a powerful method for sample efficient black box optimization. High dimensional (discrete) problems have long been identified as a key challenge for Bayesian optimization algorithms (Wang et al.,, 2013; Snoek et al.,, 2012) given that they tend to scale poorly with both dataset size and dimensionality of the input.
High-dimensional BO has been the focus of an entire research field (see Fig. 1), in which methods are extended to address the curse of dimensionality and its consequences (Binois and Wycoff,, 2022; Santoni et al.,, 2023). Within this setting, discrete sequence optimization has received particular focus, due to its applicability in the optimization of molecules and proteins. However, prior work often focuses on sequence lengths and number of categories below the hundreds (see Fig. 2), making it difficult for practitioners to judge expected performance on real-world problems in these domains. We contribute (i) a survey of the field while focusing on the real-world applications of high-dimensional discrete sequences, (ii) a benchmark of several optimizers on established black boxes, and (iii) an open source, unified interface: poli and poli-baselines.111https://github.com/MachineLearningLifeScience/{poli, poli-baselines}
@@ 2 Preliminaries
Bayesian optimization requires a surrogate model and an acquisition function (Garnett,, 2023). Given both, the objective function is sequentially optimized by fitting a model to the given observations and numerically optimizing the acquisition function with respect to the model to select the next configuration for evaluation. Frequently, the model is a Gaussian process (GP, Rasmussen and Williams, (2006)), and popular choices for the acquisition function are Expected Improvement (Jones et al.,, 1998; Garnett,, 2023) and the Upper Confidence Bound (Srinivas et al.,, 2012). A GP allows to express a prior belief over functions. Formally, it is a collection of random variables, such that every finite subset follows a multivariate normal distribution, described by a mean function , and a positive definite covariance function (kernel) (Rasmussen and Williams,, 2006, p. 13). Assuming that observations of the function are distorted by Gaussian noise, the posterior over the function conditioned on these observations is again Gaussian. The prediction equations have a closed form and can be evaluated in time where is the number of observations.
@@
Most HDBO methods are tested on toy examples, hyperparameter tuning, or reinforcement learning tasks (Binois and Wycoff,, 2022; Penubothula et al.,, 2021). We focus on discrete sequence optimization, which has several applications beyond the usual examples (e.g. MaxSAT, or PestControl) (Papenmeier et al.,, 2024), and is key in applications to biology and bioinformatics (Gómez-Bombarelli et al.,, 2018; Stanton et al.,, 2022; Gruver et al.,, 2023). Drug design and protein engineering can be thought of as sequence optimization problems, if we consider the SMILES/SELFIES representation of small molecules (Weininger,, 1988; Krenn et al.,, 2020), or the amino acid sequence representation of proteins (Needleman and Wunsch,, 1970).
@@
Binois and Wycoff, (2022) initially surveyed the field of high-dimensional GPs, focusing on applications to BO, and proposed a taxonomy of structural assumptions for GPs that includes variable selection, additive models, linear, and non-linear embeddings. This work has since been updated by Wang et al., (2023) and Santoni et al., (2023). The latter presents an empirical study on the continuous, toy-problem setting up to 60 dimensions and refines the taxonomy (Binois and Wycoff,, 2022) into five categories, separating trust regions from the rest. Griffiths et al., (2023) compare different kernel functions used with binary vector representations of molecules and for the same application Kristiadi et al., (2024) study the use of different large language models. Our work is most similar to Dreczkowski et al., (2023)’s comprehensive overview of discrete BO (MCBO), and Gao et al., (2022)’s benchmark of small molecule optimization. In both, HDBO is not in focus.
@@ 3 A taxonomy of high-dimensional Bayesian Optimization
We describe the field of high dimensional BO and the large number of related publications through a refined taxonomy building on previous work, discussing variable selection, additive models, trust regions, linear embeddings, non-linear embeddings, gradient information, structured spaces, and others in turn. While encompassing taxonomies over fields may initially appear ill-advised (Wilkins,, 1668, pp.22), we highlight commonalities in strategies that give structure to the HDBO problem-space.
We expand previous surveys (Binois and Wycoff,, 2022; Santoni et al.,, 2023) and identify a finer taxonomy of seven method groups and new families of structured spaces (i.e. methods that work directly on mixed representations, or Riemannian manifolds, previously categorized as non-linear embeddings), and methods that rely on predicted gradient information. This new separation emphasizes the heterogeneous nature of discrete solvers: some optimizers work directly on discrete space (structured spaces), while others optimize using latent representations (non-linear embeddings); gradient-based methods are separated to show alternatives when first-order information is available or modelable. Fig. 1 presents a timeline of HDBO methods, split into these families, and all methods are detailed in supplementary Table LABEL:tab:appendix:full_taxonomy; methods are grouped according to their most dominant feature.
@@ 3.1 Variable selection
To solve a high-dimensional problem, one approach is to focus on a subset of variables of high interest.222Under the assumption that there exists an axis-aligned lower-dimensional active subspace. One selects the variables either by using domain expertise, or by Automatic Relevance Detection (ARD) (Rasmussen and Williams,, 2006, pp.106-107) i.e. large lengthscales indicate independence under the covariance matrix for GPs. Examples of this approach include Hierarchical Diagonal Sampling (HDS) (Chen et al.,, 2012) and the Dimension Scheduling Algorithm (DSA) (Ulmasov et al.,, 2016). The former determines the active variables by a binary tree of subsets of , and fits GPs in lower-dimensional projections. DSA constructs a probability distribution by the principal directions of the training inputs and subsamples the dimensions accordingly. In contrast Li et al., (2018) randomly sample subsets of active dimensions.
@@ 3.2 Additive models
Han et al., (2021) select the decomposition which maximizes the marginal likelihood from a collection of randomly sampled decompositions, updating it every certain number of iterations. Alternatives include: leveraging a generalization based on restricted projections (Li et al.,, 2016), discovering the additive structure using model selection and Markov Chain Monte Carlo (Gardner et al.,, 2017), considering overlapping groups (Rolland et al.,, 2018), ensembles of Mondrian space-tiling trees (Wang et al.,, 2018), or use random tree-based decompositions (Ziomek and Bou Ammar,, 2023).
@@ 3.4 Linear embeddings
Instead of optimizing directly in input space , several methods rely on optimizing in a lower-dimensional space , which is linearly embedded into data space using a linear transformation (Wang et al.,, 2016). The matrix can be either selected at random (Wang et al.,, 2016; Qian et al.,, 2016), computed as a low-rank approximation of the input data matrix (Djolonga et al.,, 2013; Zhang et al.,, 2019; Raponi et al.,, 2020), constructed using gradient information and active subspaces (Palar and Shimoyama,, 2017; Wycoff et al.,, 2021), or through the minimization of variance estimates (Hu et al.,, 2024).
@@ 3.5 Non-linear embeddings
Gómez-Bombarelli et al., (2018) pioneered latent space optimization (LSBO) by learning a latent space of small molecules through their SMILES representation using a Variational Autoencoder (VAE, Kingma and Welling, (2014); Rezende et al., (2014)), and optimizing metrics such as the qualitative estimate of druglikeness (QED) therein. Several approaches have followed, including usage of a-priori given labelled data (Eissman et al.,, 2018) or decoder uncertainty (Notin et al.,, 2021), smart retraining schemes that focus on promising points (Tripp et al.,, 2020), metric-learning approaches that match promising points together (Grosnit et al.,, 2021), constraining the latent space (Griffiths and Hernández-Lobato,, 2020), latent spaces mapping to graphs (Kusner et al.,, 2017; Jin et al.,, 2018) and jointly learning the surrogate model and the latent representation (Maus et al.,, 2022; Lee et al.,, 2023; Chen et al.,, 2024; Kong et al.,, 2024). Stanton et al., (2022) take this further by learning multiple representations: one shared and required for both the decoder and surrogate, and one discriminative encoding as input for a GP used in the acquisition function. A prerequisite for these methods is a large dataset of unsupervised inputs, which may not be available in all applications. The methods that rely on training both the representation and the regression at the same time need supervised labels, which may be potentially unavailable. Optimization in embedding spaces greatly increases the complexity of problems that can be tackled, making it an appealing alternative for discrete sequence optimization in real-world tasks (see Fig. 2).
@@ 3.7 Structured spaces
Some applications work over structured spaces. For example, the angles of robot arms and protein backbones map to Riemannian manifolds (Jaquier et al.,, 2020; Penner,, 2022), and input spaces might also contain mixed variables (i.e. products of real and categorical spaces). To compute non-linear embeddings (see Sec. 3.5) followed by standard Bayesian optimization (or small variations thereof) can allow us to work over such spaces. Jaquier et al., (2020) use kernels defined on Riemannian manifolds (Feragen et al.,, 2015; Borovitskiy et al.,, 2020) and optimize the acquisition function using tools from Riemannian optimization (Boumal,, 2023). The authors expand their framework to high-dimensional manifolds by projecting to lower-dimensional submanifolds, which is roughly the equivalent to linear embeddings in the Riemannian settings (Jaquier and Rozo,, 2020).
@@ 4 Benchmarking the performance of HDBO methods
Practitioners that decide on what Bayesian optimization algorithm to use for their application will face several challenges. While surveying the field, we noticed two key discrepancies in the reported experimental set-ups: (i) the initialization varies from as low as none randomly/SOBOL sampled points to over , (ii) evaluation budgets also vary for the same types of tasks. Fig. 3 visualizes these different experimental set-ups as swarmplots. Moreover, our survey covered code availability. The state-of-the-art is being pulled by workhorses, which have democratized access to GP and BO implementations: GPyTorch (Gardner et al.,, 2018) and BoTorch (Balandat et al.,, 2020), and GPFlow (Matthews et al.,, 2017; van der Wilk et al.,, 2020) and Trieste (Picheny et al.,, 2023). These libraries are highly useful and impactful, yet one can obtain cross-dependency conflicts between them especially if third-party dependencies are introduced or if very specific versions are required for solver setups. As a particular example, solvers like ProbRep cannot co-exist with Ax-based solvers like SAASBO or Hvarfner’s VanillaBO. There is a need for isolating optimizers, specifying up-to-date environments in which they can run. These issues led to the development of poli.
@@
We want to solve truly high dimensional problems that are relevant for domains like biology and chemistry. To make the outcomes comparable, we require a unified way of defining the problem which includes consistent starting points, budgets, runtime environments, relevant assets (i.e.models used for the black box), and a logging backend invoked for every oracle observation. To that end, we implement the Protein Objectives Library (poli) to provide potentially isolated black box functions. Building on open source tools, poli currently provides 35 black box tasks; besides the Practical Molecular Optimization (PMO) benchmark (Huang et al.,, 2021; Gao et al.,, 2022; Brown et al.,, 2019), it includes Dockstring (García-Ortegón et al.,, 2022) as well as other protein-related black boxes like stability and solvent accessibility (Delgado et al.,, 2019; Blaabjerg et al.,, 2023; Chapman and Chang,, 2000; Stanton et al.,, 2022).333A complete list can be found in our documentation: https://machinelearninglifescience.github.io/poli-docs/ The majority of black box functions can be queried with any string that complies with the corresponding alphabet making the oracles available for free-form optimization. This is an important distinction compared to pre-existing benchmarks that rely on pools of precompiled observations (Notin et al.,, 2023).
We further provide an interface for the solvers used for the individual optimization tasks: poli-baselines. Consistent, stable (and up to date) environments of individual optimizers can be found therein, as well as a standardized way to query them and solve the problems raised in the previous section. These environments and optimizers are tested weekly through GitHub actions, guaranteeing their usability. An example of how poli and poli-baselines work is provided above. A problem containing a black box, an initial input, and potentially a data package is created through problem factories/a create method. A solver takes as input a black box and optionally supervised data, and uses the solve method to run the optimization for a given number of iterations. Such an interface can accommodate any optimizer that provides a method for running the optimization for a given amount of steps, as well as optimizers that do not accept custom initializations (e.g. Bounce or BAxUS). Sec. A.4 provides an introduction to this software’s technical details, including a description of the logging logic though observers that keep track of every black box call.
[…]
DirectedEvolution | |||||
HillClimbing | |||||
CMAES | |||||
GeneticAlgorithm | |||||
Hvarfner’s VanillaBO | |||||
RandomLineBO | |||||
SAASBO | |||||
Turbo | |||||
BAxUS | |||||
Bounce | |||||
ProbRep |
4.2 Benchmarking HDBO in discrete sequences
Using the black boxes provided in poli, as well as the solvers provided in poli-baselines, we benchmark the performance of high-dimensional Bayesian optimization solvers on discrete sequences. Such optimization can take place either at the sequence level (as solvers in the Structured Spaces do), or in a continuous version via either one-hot representations or learned latent spaces. Benchmarks on both fronts are presented in this paper: from sequence design tasks of varying complexity (mimicking protein engineering) in one-hot/sequence space using Ehrlich functions (Stanton et al.,, 2024), to latent space optimization of small molecules on the Practical Molecular Optimization (PMO) benchmark (Gao et al.,, 2022; Huang et al.,, 2021; Brown et al.,, 2019).444We emphasize that the results provided in this paper are continuously updated on our project’s website.
4.2.1 Sequence design problems of varying complexity
Ehrlich functions (Stanton et al.,, 2024) are closed-form procedurally-generated oracles in which a certain collection of motifs needs to be satisfied in a sequence of a pre-selected length. The oracle’s score is between 0 and 1, and determined by how much of the individual motifs are satisfied. The number of motifs and their length are hyperparameters specified by the user. Sec. A.5.1 provides a detailed introduction. In particular, we test on 4 different configurations: a PestControl equivalent555PestControl is a black box whose input are 25 categorical variables, each of which has 5 potential categories. It is a commonly tested black box in the Structured Spaces literature. with an alphabet size of 5 and a sequence length of 25 (i.e. one motif of length 25). Further, 3 different configurations imitating protein design using the alphabet of 20 amino acids and sequence lengths of 5, 15, and 64 are tested. The motif lengths and number of motifs are (4, 1), (7, 2) and (10, 4) respectively. We initialize PestControlEquiv, and the two small Ehrlich problems with 10 supervised samples, and the latter with 1000, and all methods have an evaluation budget of 300.
We optimize these oracles with representatives from the taxonomy that are frequently tested in the HDBO literature. We select Hvarfner’s VanillaBO, RandomLineBO, Turbo, BAxUS, SAASBO, Bounce, and ProbRep, including also baselines like DirectedEvolution (i.e. greedily mutating the incumbent in sequence space), HillClimbing (which explores the input space by taking random Gaussian steps), and evolutionary algorithms/strategies like GeneticAlgorithm and CMA-ES. Notice that, due to the nature of their implementations, BAxUS and Bounce are not initialized with the same supervised data (but rather according to their implementation: SOBOL sampling in a linear subspace). In this scenario, the continuous solvers are optimizing over one-hot inputs, and the discrete ones work over sequence space. The next section explores latent space optimization.
Table 1 shows the mean best value achieved in the aforementioned 4 problems with one standard deviation for 5 seeds. Problems of low complexity (i.e. PestControlEquiv and Ehrlich with seq. length 5) are readily solvable using one-hot BO, with Bounce solving PestControlEquiv in all 5 replications. Such problems are also easily solvable by naïve baselines like DirectedEvolution. For more complex problems, BO optimizers perform equally or worse than greedy baselines. We hypothesize that naïve baselines will perform worse on black boxes with higher-order effects. Note that SAASBO and ProbRep ran out of memory when instantiating the largest problem.
4.2.2 Benchmarking HDBO on PMO
Hill | Genetic | CMAES | Hvarfner’s | Random | SAASBO | BAxUS | Turbo | Bounce | ProbRep | |
Oracle | Climbing | Algorithm | VanillaBO | LineBO | ||||||
albuterol_similarity | ||||||||||
amlodipine_mpo | ||||||||||
celecoxib_rediscovery | ||||||||||
deco_hop | ||||||||||
drd2_docking | ||||||||||
fexofenadine_mpo | ||||||||||
gsk3_beta | ||||||||||
isomer_c7h8n2o2 | ||||||||||
isomer_c9h10n2o2pf2cl | ||||||||||
jnk3 | ||||||||||
median_1 | ||||||||||
median_2 | ||||||||||
mestranol_similarity | ||||||||||
osimetrinib_mpo | ||||||||||
perindopril_mpo | ||||||||||
ranolazine_mpo | ||||||||||
rdkit_logp | ||||||||||
rdkit_qed | ||||||||||
sa_tdc | ||||||||||
scaffold_hop | ||||||||||
sitagliptin_mpo | ||||||||||
thiothixene_rediscovery | ||||||||||
troglitazone_rediscovery | ||||||||||
valsartan_smarts | ||||||||||
zaleplon_mpo | ||||||||||